What is inverse of i(iota) ?

Grade:12th pass

5 Answers

Ashwini
9 years ago
I dont know.
Ashwini
9 years ago
How would I know?
Shivam Chopra
9 years ago
It is -i.
It is because in verse of i = 1/i.
Multiplying & dividing by i, we get
1/i = i/i2
Also, i2 = -1
So, 1/i = i/-1 = -i
Please approve it if you like.
rahul
8 years ago
here we just have to rationalise  1/i
this will give us the result -i
 
hit approve if useful
abhinaipadachuru
8 years ago
It is -1 because inverse means1/i=-i,divide numarater and denominatot by i the in numarater you have i and in denominator i square.isquare=-1.So the answer is -i.
In future try to look at some universities which provide excellent placement opportunities,top infrastructure.
The universities are
SSN College of Engineering, Chennai
Lovely Professional University.
SJ College of Engineering, Mysore
SASTRA University, Thanjavur.
For more information visit websites.